Suppose Paul's utility depends on the amount of time spent playing on the Internet (x) and the amount of time playing video games (y), and his utility function is U(x,y) = xy/(x+y). He has 20 hours of free time to spend on these two activities each week, and his goal is to maximize his utility. What is the optimal amount of time spent surfing the Internet and playing video games each week?
Answer #1

The utility function is h + -= LU , while the budget constraint would be r+y = 20 . The Lagrangian function to maximize the utility subjected to the constraint would be as L= 19 1 + y + (20-r-y) . The FOCs would be as below.

\frac{\partial L}{\partial x} = 0 or V 9 + (20-r-y)) = 0 2 + y or 19 r+ y a +w21(-1) = 0 or y(1+y) (x + y)2 ( ry + y)2 or ry + y (x + y)2 ry (+y) or (もっ .

\frac{\partial L}{\partial y} = 0 or a ry +(20-1-y)) = 0 or LU r+y (r + y)2 12 + (-1)=0 or r(r + y) (x + y)2 ry + y)2 ( or r+ rg ( + y )2 ᎢᏴ {r +1 )2- or (r - 2 =A .

\frac{\partial L}{\partial \lambda} = 0 or 0 0 + 2x (20-r-y)) = 0 + y or 20-r- y) = 0 or r+y = 20 .

Comparing the first two FOCs, we have (r + y)2 (r + y)2 or y^2 = x^2 or y = \pm x , and since x and y can not be negative, we have y = x as the required condition for utility maximizing combination of x and y.

Putting this in the constraint, we have r+ (r) = 20 or 2x = 20 or x^* = 10 , and since y = x , we have y^* = x^* or y^* = 10 .

Hence, the optimal amount if time spent will be 10 hours of internet surfing and 10 hours of gaming.